# MIMO Detection with Spatial Sigma-Delta ADCs: A Variational Bayesian Approach

### v1

@@ Abstract
The spatial Sigma-Delta ($\Sigma\Delta$) architecture can be leveraged to reduce the quantization noise and enhance the effective resolution of few-bit analog-to-digital converters (ADCs) at certain spatial frequencies of interest. Utilizing the variational Bayesian (VB) inference framework, this paper develops novel data detection algorithms tailored for massive multiple-input multiple-output (MIMO) systems with few-bit $\Sigma\Delta$ ADCs and angular channel models, where uplink signals are confined to a specific angular sector. We start by modeling the corresponding Bayesian networks for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ receivers. Next, we propose an iterative algorithm, referred to as Sigma-Delta variational Bayes (SD-VB), for MIMO detection, offering low-complexity updates through closed-form expressions of the variational densities of the latent variables. Simulation results show that the proposed $2^{\rm nd}$-order SD-VB algorithm delivers the best symbol error rate (SER) performance while maintaining the same computational complexity as in unquantized systems, matched-filtering VB with conventional quantization, and linear minimum mean-squared error (LMMSE) methods. Moreover, the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms achieve their lowest SER at an antenna separation of one-fourth wavelength for a fixed number of antenna elements. The effects of the steering angle of the $\Sigma\Delta$ architecture, the number of ADC resolution bits, and the number of antennas and users are also extensively analyzed.
Index Terms:
Few-bit quantization, MIMO detection, spatial Sigma-Delta ADCs, variational Bayesian inference.
@@ I Introduction
Beyond-5G wireless systems will require substantial bandwidth in both the millimeter (mmWave) and terahertz (THz) bands to deliver high data throughput [2]. Signals at these frequencies, however, are hindered by low penetration capabilities and high propagation loss, which restrict their practical communication range [3]. Massive multiple-input multiple-output (MIMO) arrays have been used to compensate for the propagation loss while simultaneously achieving high capacity through spatial multiplexing [4]. However, exploiting the full benefits of beamforming and multiplexing in massive MIMO can be challenging due to the need for dedicated high-resolution analog-to-digital converters (ADCs)/digital-to-analog converters (DACs) for each antenna element [5, 6]. This results in high hardware complexity and increased power consumption, especially with larger bandwidths and sampling rates [7]. To address these concerns, the use of low-resolution ADCs with, e.g., 11113333 bits of precision, has emerged as an energy-efficient and low-complexity solution for massive MIMO systems [8, 9, 10].
Low-resolution ADCs provide low-complexity and low-power designs for massive MIMO systems, but present challenges due to their non-linear nature [11]. The impact of low-resolution ADCs on MIMO systems has been explored, including analyses of the resulting capacity limits and achievable rates [12, 13, 14, 11, 15]. These studies demonstrate that low-resolution ADCs degrade system performance metrics such as achievable rate and symbol error rate (SER), especially at medium-to-high signal-to-noise ratios (SNRs), but this negative impact can be mitigated by increasing the number of antennas. This observation indicates that massive MIMO can be effectively operated with few-bit ADCs. However, attaining a good tradeoff between the system performance and the number of quantization bits requires advanced signal processing algorithms and architectural innovations tailored for the specific characteristics of the quantized signals [16, 10].
@@ I-A Related Works
Communications in the mmWave and THz bands typically require large antenna arrays at the base stations (BS). Here, the implementation of one-bit ADCs for every antenna element can reduce radio-frequency complexity, cost, and power consumption [22, 23]. In [24], a likelihood function learning method for data detection was proposed with one-bit ADCs using an approach based on reinforcement learning. In [16, 25], a support vector machine (SVM)-based channel estimation and data detection method was exploited for one-bit massive MIMO systems, yielding a data detection performance close to that of maximum-likelihood detection. Data detection based on supervised learning was proposed for few-bit MIMO systems in [26], whereas an approach based on deep neural networks was developed in [27]. These learning-based methods necessitate a significant training load, and the neural network must be retrained with new channel realizations [27], thereby increasing the computational complexity. To efficiently balance the performance and complexity for systems with few-bit ADCs, variational Bayes (VB) inference, inspired by machine learning, was proposed for data detection and channel estimation in [10]. The goal of VB is to find an approximation for the true posterior distribution of latent variables given the observed data [28]. VB has also been used in the context of joint channel estimation and data detection in massive MIMO [29] and orthogonal frequency-division multiplexing systems with few-bit ADCs [30]. However, there is no prior work that has considered VB for MIMO detection with $\Sigma\Delta$ ADCs.
@@
Recently, spatial $\Sigma\Delta$ ADCs in massive MIMO systems have attracted growing interest in terms of channel estimation [18, 21], performance analysis [18, 20, 31], and precoding design [32]. In [20], the spectral efficiency of $1$-bit $\Sigma\Delta$ massive MIMO was analyzed, revealing that one-bit $\Sigma\Delta$ scales down the quantization noise power proportionally to the square of the spatial oversampling rate. Spatial few-bit $\Sigma\Delta$ ADCs have also been used in massive MIMO to shape the quantization noise away from users in certain angular sectors for improved channel estimation [18]. In [31], a parallel $\Sigma\Delta$ ADC was designed for mmWave receivers to achieve high-resolution outputs; however, the bit-error-rate performance of this design with $3$-bit quantization was not significantly improved at high SNRs. $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ quantization was employed for wide-band systems in [33] without considering the angle steering mechanism. A reconfigurable intelligent surface controlled by a single-antenna base station equipped with $1^{\rm st}$-order $\Sigma\Delta$ modulation was investigated in [34]. Mutual coupling effects for phase quantization were considered for dense $\Sigma\Delta$ phased arrays in [35].
@@ I-B Motivations and Contributions
A common approach in the aforementioned works on $\Sigma\Delta$ ADCs is to exploit the Bussgang decomposition, which reformulates the nonlinear quantization as a linear function, followed for example by standard linear data detection and channel estimation methods [11, 32, 18]. However, the Bussgang decomposition is based on the assumption of Gaussian signals, which may not hold at high SNR when the channels are specular and there are few users. The VB method has been proposed as an alternative to effectively address the non-linear quantization operation. It is considered superior to linear methods, bilinear generalized approximate message passing [29], and other approaches based on deep learning [36]. However, these studies are limited to conventional few-bit ADCs. More importantly, achieving optimal MIMO detection with spatial $\Sigma\Delta$ ADCs under the maximum-a-posteriori (MAP) probability criterion may not be feasible. These observations prompt the development in this paper of novel data detection algorithms based on the VB inference framework for MIMO systems with spatial $\Sigma\Delta$ ADCs.
@@
  • We develop $1^{\rm st}$- and $2^{\rm nd}$-order Sigma-Delta Variational Bayes (SD-VB) detection algorithms, where few-bit $\Sigma\Delta$ quantizers are employed in a massive MIMO system. We model the Bayesian network for the $\Sigma\Delta$ receivers, which reveals insightful dependencies between the observed and latent variables.
  • We develop the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ ADC architectures with efficient updates based on the analysis of the variational densities. We also present an LMMSE detector for MIMO detection with $1^{\rm st}$-order $\Sigma\Delta$ quantization. The results show that the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms have the same computational complexity as the matched-filter quantized VB (MF-QVB) algorithm [10], and significantly lower complexity than the LMMSE detector.
  • We demonstrate through simulation results that the $2^{\rm nd}$-order SD-VB algorithm achieves the lowest SER compared with state-of-the-art detection algorithms such as MF-QVB and LMMSE for certain angular sectors. Moreover, the $1^{\rm st}$-order SD-VB algorithm is more efficient than $2^{\rm nd}$-order SD-VB for $1$-bit ADCs.
The rest of the paper is organized as follows. Section II introduces the system model as well as the model for few-bit $\Sigma\Delta$ ADCs. The optimal solution for the LMMSE detector based on the linearized model is derived in Section III. Section IV analyzes and discusses the VB framework for detection in a $1^{\rm st}$-order $\Sigma\Delta$ MIMO system. Then, Section V extends the VB framework to the $2^{\rm nd}$-order $\Sigma\Delta$ case. Numerical results are provided in Section VI and Section VII concludes the paper.
[…]
II System Model
@@ II-A Channel Model
We consider an uplink MIMO system consisting of $K$ single-antenna users and a BS equipped with a uniform linear array (ULA) of $N$ antennas. The received signal $\mathbf{x}$ at the BS before quantization can be expressed as
@@
where $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{N}]^{T}\in\mathbb{C}^{N\times K}$ is the uplink channel matrix, with $\mathbf{h}_{i}\in\mathbb{C}^{K\times 1}$ denoting the channel vector from the $K$ users to the $i$-th antenna at the BS, $\mathbf{s}=[s_{1},\ldots,s_{K}]^{T}\in\mathbb{C}^{K\times 1}$ is the uplink data symbol vector, and $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},N_{0}\mathbf{I}_{N})$ represents additive Gaussian noise at the BS with variance $N_{0}$. We assume that the channel $\mathbf{H}$ is known at the BS. The symbol $s_{k}$ transmitted from user-$k$ is drawn from a complex-valued discrete constellation $\mathcal{S}$, e.g., quadrature amplitude modulation (QAM) or phase-shift keying (PSK), with $\mathbb{E}[s_{k}]=0$ and normalized such that $\mathbb{E}[|s_{k}|^{2}]=1$. The prior distribution of $s_{k}$ can be expressed as
@@
$\bar{\mathbf{h}}_{k}=\sqrt{\frac{\beta_{k}}{L}}\mathbf{A}_{k}\mathbf{g}_{k},$ (3)
where $\bar{\mathbf{h}}_{k}$ is the $k$th column of $\mathbf{H}$ and the uplink channel from user-$k$, the columns of $\mathbf{A}_{k}\in\mathbb{C}^{N\times L}$ represent the array response for $L$ propagation paths, $\mathbf{g}_{k}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{L})$ represents the small-scale fading, and $\beta_{k}$ models the geometric attenuation and slow fading. We assume that $\mathbf{A}_{k}$ is a full rank matrix whose $\ell$-th column is the array steering vector corresponding to the angle of arrival (AoA) $\theta_{k\ell}$ of the $\ell$-th path, as given by
@@
where $d$ is the inter-antenna spacing of the ULA and $\lambda$ denotes the wavelength. For simplicity, we denote $\omega_{{k\ell}}=2\pi\frac{d}{\lambda}\sin\theta_{k\ell}$ as the spatial frequency of the $\ell$-th path from user-$k$. We assume that the AoAs for all users are situated within a specific angular sector $\mathcal{S}_{\theta_{0}}\triangleq[\theta_{0}-\Theta/2,\theta_{0}+\Theta/2]$, where $\theta_{0}$ is the center angle of the sector and $\Theta$ is the azimuth angular spread. This constraint frequently arises in practice because cell sectoring limits service to users within a specific angular sector of the BS coverage. Additionally, the high frequency and directional nature of mmWave signals produce narrow beam patterns and substantial attenuation of distant reflections, making nearby reflections of users more dominant and resulting in narrower angular spreads for users.
The received analog signals are then quantized using few-bit ADCs. The mathematical models for both conventional and $\Sigma\Delta$ quantizers in these few-bit ADCs will be detailed in the following subsection.

### v2

@@ Abstract
The spatial Sigma-Delta ($\Sigma\Delta$) architecture can be leveraged to reduce the quantization noise and enhance the effective resolution of few-bit analog-to-digital converters (ADCs) at certain spatial frequencies of interest. Utilizing the variational Bayesian (VB) inference framework, this paper develops novel data detection algorithms tailored for massive multiple-input multiple-output (MIMO) systems with few-bit $\Sigma\Delta$ ADCs and angular channel models, where uplink signals are confined to a specific angular sector. We start by modeling the corresponding Bayesian networks for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ receivers. Next, we propose an iterative algorithm, referred to as Sigma-Delta variational Bayes (SD-VB), for MIMO detection, offering low-complexity updates through closed-form expressions of the variational densities of the latent variables. We also study the impact of mutual coupling on the performance of the proposed SD-VB algorithms when the antenna spacing is reduced. Simulation results show that the proposed $2^{\rm nd}$-order SD-VB algorithm delivers the best symbol error rate (SER) performance while maintaining the same computational complexity as in unquantized systems, matched-filtering VB with conventional quantization, and linear minimum mean-squared error (LMMSE) methods. Moreover, the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms achieve their lowest SER at an antenna separation of one-fourth wavelength for a fixed number of antenna elements. The effects of mutual coupling, the steering angle of the $\Sigma\Delta$ architecture, the number of ADC resolution bits, and the number of antennas and users are also extensively analyzed.
@@ I Introduction
Beyond-5G wireless systems will require substantial bandwidth in both the millimeter (mmWave) and terahertz (THz) bands to deliver high data throughput [2]. Signals at these frequencies, however, are hindered by low penetration capabilities and high propagation loss, which restrict their practical communication range [3]. Massive multiple-input multiple-output (MIMO) arrays have been used to compensate for the propagation loss while simultaneously achieving high capacity through spatial multiplexing [4]. However, exploiting the full benefits of beamforming and multiplexing in massive MIMO can be challenging due to the need for dedicated high-resolution analog-to-digital converters (ADCs)/digital-to-analog converters (DACs) for each antenna element [5, 6]. This results in high hardware complexity and increased power consumption, especially with larger bandwidths and sampling rates [7]. To address these concerns, the use of low-resolution ADCs with, e.g., 1133 bits of precision, has emerged as an energy-efficient and low-complexity solution for massive MIMO systems [8, 9, 10].
Low-resolution ADCs provide low-complexity and low-power designs for massive MIMO systems, but present challenges due to their non-linear nature [11]. The impact of low-resolution ADCs on MIMO systems has been explored, including analyses of the resulting capacity limits and achievable rates [12, 13, 14, 11, 15]. These studies show that low-resolution ADCs degrade system performance metrics such as achievable rate and symbol error rate (SER), especially at medium-to-high signal-to-noise ratios (SNRs), but this negative impact can be mitigated by increasing the number of antennas. This observation indicates that massive MIMO can be effectively operated with few-bit ADCs. However, attaining a good tradeoff between the system performance and the number of quantization bits requires advanced signal processing algorithms and architectural innovations tailored for the specific characteristics of the quantized signals [16, 10].
@@ I-A Related Works
Communications in the mmWave and THz bands typically require large antenna arrays at the base stations (BS). Here, the implementation of one-bit ADCs for every antenna element can reduce radio-frequency complexity, cost, and power consumption [22, 23]. In [24], a likelihood function learning method for data detection was proposed with one-bit ADCs using an approach based on reinforcement learning. In [16, 25], a support vector machine (SVM)-based channel estimation and data detection method was exploited for one-bit massive MIMO systems, yielding a data detection performance close to that of maximum-likelihood detection. Data detection based on supervised learning was proposed for few-bit MIMO systems in [26], whereas an approach based on deep neural networks was developed in [27]. These learning-based methods necessitate a significant training load, and the neural network must be retrained with new channel realizations [27], thereby increasing the computational complexity. To reduce computational complexity, zero-forcing-based data detection was studied in uplink massive MIMO systems with few-bit ADCs in [28].
Generalized approximate message passing (GAMP)-based algorithms, such as generalized orthogonal approximate message passing (GOAMP) [29] and generalized memory approximate message passing (GMAMP) [30], have also been proposed for generalized linear models with few-bit ADCs. These algorithms can be applied to any discrete input signals and achieve Bayesian optimality when the transformation matrix is high-dimensional and is unitarily invariant. A state evolution framework was used in these works to analyze the performance of GAMP-based algorithms in the large-system limit. However, for mmWave channel models with non-unitarily invariant channel structures and highly non-linear, non-separable spatial $\Sigma\Delta$ quantization, there is no guarantee that GOAMP and GMAMP will converge or achieve Bayesian optimality, motivating the need for alternative detection approaches. The use of analog comparator networks prior to 1-bit quantization was introduced in [31] to effectively create virtual channels whose outputs can enhance the performance of MIMO receivers. To further reduce the effect of non-linear quantization, the authors of [32] studied the integration of pre-ADC comparators and deep learning detection in orbital angular momentum systems with 1-bit spatial oversampling.
To efficiently balance the performance and complexity for systems with few-bit ADCs, variational Bayesian (VB) inference, inspired by machine learning, was proposed for data detection and channel estimation in [10]. The goal of VB is to find an approximation for the true posterior distribution of latent variables given the observed data [33]. In [34], the VB method was applied for gridless channel estimation in antenna array systems with low-resolution ADCs.
@@
Recently, spatial $\Sigma\Delta$ ADCs in massive MIMO systems have attracted growing interest in terms of channel estimation [18, 21], performance analysis [18, 20, 35], and precoding design [36]. In [20], the spectral efficiency of $1$-bit $\Sigma\Delta$ massive MIMO was analyzed, revealing that one-bit $\Sigma\Delta$ scales down the quantization noise power proportionally to the square of the spatial oversampling rate. Spatial few-bit $\Sigma\Delta$ ADCs have also been used in massive MIMO to shape the quantization noise away from users in certain angular sectors for improved channel estimation [18]. In [35], a parallel $\Sigma\Delta$ ADC was designed for mmWave receivers to achieve high-resolution outputs; however, the bit-error-rate performance of this design with $3$-bit quantization was not significantly improved at high SNRs. $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ quantization was employed for wide-band systems in [37] without considering the angle steering mechanism. A reconfigurable intelligent surface controlled by a single-antenna base station equipped with $1^{\rm st}$-order $\Sigma\Delta$ modulation was investigated in [38]. Mutual coupling effects for phase quantization were considered for dense $\Sigma\Delta$ phased arrays in [39]. A VB algorithm for optimal detection in MIMO systems using spatial $\Sigma\Delta$ ADCs was recently developed in [1]. This approach leverages the noise-shaping properties of the $1^{\rm st}$-order $\Sigma\Delta$ architecture to improve detection performance in angularly constrained scenarios. However, the study remains limited to $1^{\rm st}$-order $\Sigma\Delta$ ADCs, which do not fully capitalize on the potential performance gains offered by higher-order $\Sigma\Delta$ architectures.
@@ I-B Motivations and Contributions
A common approach in the aforementioned works on $\Sigma\Delta$ ADCs is to exploit the Bussgang decomposition, which reformulates the nonlinear quantization as a linear function, followed for example by standard linear data detection and channel estimation methods [11, 36, 18]. However, the Bussgang decomposition relies on the assumption of Gaussian signals, which results in a diagonal Bussgang gain matrix. At high SNR with specular channels [40] and relatively few users, the Gaussian assumption does not hold and the Bussgang gain matrix becomes non-diagonal, leading to correlated distortion terms across users or spatial dimensions [41, 42, 18]. For example, this effect is manifested when the nonlinear quantization generates harmonics for sparse specular inputs, as observed in [43].
The VB method has been proposed as an alternative to the Bussgang approach to effectively address the non-linear quantization operation. It is considered superior to linear methods, bilinear GAMP [44], and other approaches based on deep learning [45]. However, these studies are limited to conventional few-bit ADCs. More importantly, achieving optimal MIMO detection with spatial $\Sigma\Delta$ ADCs under the maximum-a-posteriori (MAP) probability criterion may not be feasible due to the complex noise-shaping effects that disrupt the posterior distribution. Building on our initial exploration of MIMO detection with $1^{\rm st}$-order $\Sigma\Delta$ ADCs in [1], this paper develops a comprehensive VB framework for MIMO detection that incorporates both $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ architectures. This framework addresses key limitations of prior VB studies, such as their focus on conventional ADCs and lack of support for higher-order noise shaping. Moreover, we consider a narrowband channel model typical of mmWave systems, where the signal bandwidth is smaller than the coherence bandwidth. This model is relevant to applications such as short-range indoor communications [46], mmWave MIMO integrated sensing and communication systems, and low-mobility IoT networks [47], where LoS paths or narrow angular sectors dominate over frequency selectivity. The VB inference is used to develop novel detection algorithms tailored for such scenarios.
@@
  • We develop $1^{\rm st}$- and $2^{\rm nd}$-order Sigma-Delta Variational Bayes (SD-VB) detection algorithms for massive MIMO systems with few-bit $\Sigma\Delta$ quantizers. To support this, we employ a Bayesian network to model the $\Sigma\Delta$ receiver, capturing insightful dependencies between the observed and latent variables for effective inference.
  • We develop the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ ADC architectures with efficient updates based on the analysis of the variational densities. Considering mutual coupling (MC) in $\Sigma\Delta$ architectures, we also present an LMMSE detector for MIMO detection with $1^{\rm st}$-order $\Sigma\Delta$ quantization [20]. The results show that the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms have the same computational complexity as the matched-filter quantized VB (MF-QVB) algorithm [10], and significantly lower complexity than the LMMSE detector.
  • We demonstrate through simulation results that the $2^{\rm nd}$-order SD-VB algorithm achieves the lowest SER compared with state-of-the-art detection algorithms such as MF-QVB and LMMSE for certain angular sectors. We also examine the impact of mutual coupling on the detection performance of all considered algorithms. Moreover, the $1^{\rm st}$-order SD-VB algorithm is more efficient than $2^{\rm nd}$-order SD-VB for $1$-bit ADCs.
The rest of the paper is organized as follows. Section II introduces the system model as well as the model for few-bit $\Sigma\Delta$ ADCs. The optimal solution for the LMMSE detector based on the linearized model is derived in Section II-C. Section III analyzes and discusses the VB framework for detection in a $1^{\rm st}$-order $\Sigma\Delta$ MIMO system. Then, Section IV extends the VB framework to the $2^{\rm nd}$-order $\Sigma\Delta$ case. Numerical results are provided in Section V and Section VI concludes the paper.
Notation: Boldface lowercase and boldface uppercase variables denote vectors and matrices, respectively. The $i$-th element of a vector $\mathbf{x}$ is represented by $\mathbf{x}]_{i}$ and the $(i,j)$-th element of a matrix $\mathbf{X}$ is denoted by $\mathbf{X}]_{i,j}$. The symbols $\mathbb{C}$ and $\mathbb{R}$ stand for the sets of complex and real numbers, respectively. The $L_{2}$-norm and the absolute value are indicated by $\|\cdot\|$ and $|\cdot|$, respectively. Real and imaginary parts are denoted by $\mathfrak{R}\{\cdot\}$ and $\mathfrak{I}\{\cdot\}$, respectively, with $j=\sqrt{-1}$. The identity matrix is denoted by $\mathbf{I}$, the trace operator by $\mathrm{Tr}$($\cdot$), and the expectation operator by $\mathbb{E}\{\cdot\}$. The transpose, complex conjugate, and complex conjugate transpose operators are denoted by $(\cdot)^{T}$, $(\cdot)^{*}$, and $(\cdot)^{H}$, respectively. The symbols $\propto$ and $\sim$ stand for “is proportional to” and “distributed according to”, respectively. The prior distribution function and variational distribution function of a random variable $x$ are denoted as $p(x)$ and $q(x)$, respectively. The notation $\mathcal{CN}(\bm{\mu},\bm{\Sigma})$ stands for a Gaussian random vector with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, whereas $\mathcal{CN}(\mathbf{x};\bm{\mu},\bm{\Sigma})=\frac{1}{\pi^{K}|\bm{\Sigma}|}\,\mathrm{exp}\big(-(\mathbf{x}-\bm{\mu})^{H}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\big)$ denotes the probability density function of $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\bm{\Sigma})$. The bracket $\langle.\rangle$ denotes the expectation with respect to all latent variables except the one under consideration. The cosine and sine integrals are denoted as $\mathrm{Ci}(x)\triangleq\gamma+\mathrm{log}(x)+\int_{0}^{x}{\frac{\mathrm{cos}(t)-1}{t}dt}$, and $\mathrm{Si}(x)\triangleq\int_{0}^{x}{\frac{\mathrm{sin}(t)}{t}dt}$, where $\gamma$ is the Euler-Mascheroni constant.
II System Model and Prior Art
@@ II-A Channel Model
Figure 1: Diagram of the considered system model where $K$ single-antenna users are transmitting to an NN-antenna base station employing $\Sigma\Delta$ quantization.
TABLE I: List of main notations.
Symbol Description
$K$ Number of users
NN Number of antennas at the BS
$\mathbf{x}$ Received signal vector before quantization
$\mathbf{H}$ Uplink channel matrix, $\mathbb{C}^{N\times K}$
$\mathbf{s}$ Uplink data symbol vector, $\mathbb{C}^{K\times 1}$
$\mathbf{n}$ Additive Gaussian noise, $\mathcal{CN}(\mathbf{0},N_{0}\mathbf{I}_{N})$
$N_{0}$ Noise variance
$\mathcal{S}$ Discrete constellation set (e.g., QAM, PSK)
$L$ Number of propagation paths per user
$\mathbf{A}_{k}$ Array response matrix for user $k$, $\mathbb{C}^{N\times L}$
$\mathbf{g}_{k}$ Small-scale fading vector, $\mathcal{CN}(\mathbf{0},\mathbf{I}_{L})$
$\beta_{k}$ Geometric attenuation and slow fading
$\theta_{k\ell}$ AoA of the $\ell$-th path from user $k$
$\omega_{k\ell}$ Spatial frequency, $2\pi\frac{d}{\lambda}\sin\theta_{k\ell}$
$d$ Inter-antenna spacing of the ULA
$\lambda$ Wavelength
$\mathcal{S}_{\theta_{0}}$ Angular sector, $\theta_{0}-\Theta/2,\theta_{0}+\Theta/2]$
$\theta_{0}$ Center angle of the angular sector
$\Theta$ Azimuth angular spread
$\mathbf{T}$ Mutual coupling matrix, $\mathbb{C}^{N\times N}$
$\mathbf{R}$ Covariance matrix of colored noise
$b$ Number of bits in ADC quantization
$\mathcal{Q}_{b}(\cdot)$ $b$-bit quantization function
$\Lambda$ Quantization step size
$d_{m}$ Quantization threshold, $m\in\{1,\ldots,2^{b}-1\}$
$r_{i}$ Pre-quantized signal at antenna $i$
$\phi$ Phase shift between adjacent antennas
$y_{i}$ Quantized observation at antenna $i$
$\mathbf{q}$ Effective quantization noise vector
$\mathbf{U}$ Lower triangular matrix for $\Sigma\Delta$ linearization
$\tau_{q_{i}}$ Variance of quantization noise at antenna $i$
$\mathbf{W}$ LMMSE receiver matrix
$\mathbf{\Sigma}_{\mathbf{s}}$ Covariance matrix of $\mathbf{s}$
$\mathbf{\Sigma}_{\mathbf{q}}$ Covariance matrix of $\mathbf{q}$
$\gamma$ Precision parameter, $1/N_{0}^{\text{post}}$
$\langle r_{i}\rangle$ Variational mean of $r_{i}$
$\tau_{r_{i}}$ Variational variance of $r_{i}$
We consider an uplink massive MIMO system with single-carrier spatial-division multiple access as shown in Fig. 1, where $K$ single-antenna users transmit their signals simultaneously to a BS equipped with a uniform linear array (ULA) of $N$ antennas. The received signal $\mathbf{x}$ at the BS before quantization can be expressed as
@@
where $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{N}]^{T}\in\mathbb{C}^{N\times K}$ is the uplink channel matrix, with $\mathbf{h}_{i}\in\mathbb{C}^{K\times 1}$ denoting the channel vector from the $K$ users to the $i$-th antenna at the BS, $\mathbf{s}=[s_{1},\ldots,s_{K}]^{T}\in\mathbb{C}^{K\times 1}$ is the uplink data symbol vector, and $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},N_{0}\mathbf{I}_{N})$ represents additive Gaussian noise at the BS with variance $N_{0}$. We assume that the channel $\mathbf{H}$ is known at the BS. The symbol $s_{k}$ transmitted from user $k$ is drawn from a complex-valued discrete constellation $\mathcal{S}$, e.g., quadrature amplitude modulation (QAM) or phase-shift keying (PSK), with $\mathbb{E}[s_{k}]=0$ and normalized such that $\mathbb{E}[|s_{k}|^{2}]=1$. The prior distribution of $s_{k}$ can be expressed as
@@
$\bar{\mathbf{h}}_{k}=\sqrt{{\beta_{k}}/{L}}\mathbf{A}_{k}\mathbf{g}_{k},$ (3)
where $\bar{\mathbf{h}}_{k}$ is the $k$th column of $\mathbf{H}$ and the uplink channel from user $k$, the columns of $\mathbf{A}_{k}\in\mathbb{C}^{N\times L}$ represent the array response for $L$ propagation paths, $\mathbf{g}_{k}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{L})$ represents the small-scale fading, and $\beta_{k}$ models the geometric attenuation and slow fading. We assume that $\mathbf{A}_{k}$ is a full rank matrix whose $\ell$-th column is the array steering vector corresponding to the angle of arrival (AoA) $\theta_{k\ell}$ of the $\ell$-th path, as given by
@@
where $d$ is the inter-antenna spacing of the ULA and $\lambda$ denotes the wavelength. For simplicity, we denote $\omega_{{k\ell}}=2\pi\frac{d}{\lambda}\sin\theta_{k\ell}$ as the spatial frequency of the $\ell$-th path from user $k$.
We assume that the AoAs for all users are situated within a specific angular sector $\mathcal{S}_{\theta_{0}}\triangleq[\theta_{0}-\Theta/2,\theta_{0}+\Theta/2]$, where $\theta_{0}$ is the center angle of the sector and $\Theta$ is the azimuth angular spread. This constraint frequently arises in practice because cell sectoring limits service to users within a specific angular sector of the BS coverage. Additionally, the high frequency and directional nature of mmWave signals produce narrow beam patterns and substantial attenuation of distant reflections, making nearby reflections of users more dominant and resulting in narrower angular spreads for users. Note that the channel model considered in (3) is narrowband, meaning the signal bandwidth is smaller than the coherence bandwidth. This model is relevant to applications such as short-range indoor communications [48] and low-mobility IoT networks [49], where LOS paths or narrow angular sectors dominate over frequency selectivity. This work leverages variational Bayesian inference to develop novel detection algorithms tailored for such scenarios.
II-A1 Mutual Coupling
To take mutual coupling between antennas into account, the channel vector from user $k$ to the BS can be modeled as
$\bar{\mathbf{h}}_{k}=\sqrt{{\beta_{k}}/{L}}\mathbf{T}\mathbf{A}_{k}\mathbf{g}_{k},$ (5)
where the coupling matrix $\mathbf{T}\in\mathbb{C}^{N\times N}$ is given by
$\mathbf{T}=\big(\mathbf{I}+R^{-1}\mathbf{Z}\big)^{-1}$ (6)
and $R$ denotes the low-noise amplifier (LNA) input impedance. Assuming thin half-wavelength dipoles, the elements of $\mathbf{Z}$ are expressed as [50, 18]
$[\mathbf{Z}]_{ij}=30\big(2\mathrm{Ci}(2\pi d_{ij})-\mathrm{Ci}(\xi_{ij}+\pi)-\mathrm{Ci}(\xi_{ij}-\pi)$
$\quad+j\left(-2\mathrm{Si}(2\pi d_{ij})+\mathrm{Si}(\xi_{ij}+\pi)+\mathrm{Si}(\xi_{ij}-\pi)\right)\big),~i\neq j$
$[\mathbf{Z}]_{ii}=30\bigl(\gamma+\mathrm{log}(2\pi)-\mathrm{Ci}(2\pi)+j\mathrm{Si}(2\pi)\bigr),$ (7)
where $d_{ij}$ denotes the distance between antennas $i$ and $j$ normalized by the wavelength, and $\xi_{ij}=\pi\sqrt{1+4d_{ij}^{2}}$.
Mutual coupling between antenna elements influences not only the channel characteristics but also the statistical properties of the noise at the receiver. Unlike traditional quantization schemes, the $\Sigma\Delta$ architecture leverages spatial oversampling or users confined to narrow angular regions to shape the quantization noise away from desired signal directions. However, mutual coupling introduces colored noise, denoted as $\mathbf{n}_{\rm c}\sim\mathcal{CN}(\mathbf{0},\mathbf{R})$, with covariance matrix $\mathbf{R}$. When considering $\Sigma\Delta$-based MIMO systems with less than half-wavelength-spaced antennas, it is important to consider the potential impact of mutual coupling. Based on the above electromagnetic coupling model, $\mathbf{R}$ can be expressed as [50, 18]
$\mathbf{R}=\mathbf{T}\big(\sigma_{i}^{2}(\mathbf{Z}\mathbf{Z}^{H}+R_{c}^{2}\mathbf{I}-2R_{c}\mathfrak{R}(\rho^{*}\mathbf{Z}))+4\mathrm{k}TB\mathfrak{R}(\mathbf{Z})\big)\mathbf{T}^{H},$ (8)
where $\mathbb{E}\{\mathbf{i}_{c}\mathbf{i}_{c}^{H}\}=\sigma_{i}^{2}\mathbf{I}$, $\mathbb{E}\{\mathbf{u}_{c}\mathbf{u}_{c}^{H}\}=\sigma_{u}^{2}\mathbf{I}$, $R_{c}=\frac{\sigma_{u}}{\sigma_{i}}$, $\mathbb{E}\{\mathbf{u}_{c}\mathbf{i}_{c}^{H}\}=\frac{\rho}{\sigma_{u}\sigma_{i}}\mathbf{I}$, $\mathbf{i}_{c}$ and $\mathbf{u}_{c}$ represent the equivalent noise current and voltage of the LNA, and $\mathrm{k}$, $T$, and $B$ denote the Boltzmann constant, environment temperature, and bandwidth, respectively. The received analog signals are then quantized using few-bit ADCs. The mathematical models for both conventional and $\Sigma\Delta$ quantizers in these few-bit ADCs will be presented in the following subsection.